Carbon 14 decays continously at the rate of 0.01245% per year. If 13% of the original carbon-14 remains in a specimen, estimate its age in years.
Answer by stanbon(75887)

Carbon 14 decays continously at the rate of 0.01245% per year. If 13% of the original carbon-14 remains in a specimen, estimate its age in years.---------------------------------------
A(t) = Ao^t
------
0.13Ao = Ao^t
----
0.13 = 0.999875^t
---
Take the ln of both sides:
t = ln(0.13)/ln(0.99875)
t = 1631 years
====================
Cheers,
Stan H.
====================